OpenStudy (anonymous):

Solve for x: 4/5 x + 4/3 = 2x

hero (hero):

\[\frac{4}{5}x + \frac{4}{3} = 2x\]

hero (hero):

Subtract 4/5 of x from both sides: \[\frac{4}{3} = 2x - \frac{4}{5}x\] Factor out x on the right side: \[\frac{4}{3} = \left(2 - \frac{4}{5}\right)x\] Simplify: \[\frac{4}{3} = \left(\frac{10}{5} - \frac{4}{5}\right)x\] \[\frac{4}{3} =\frac{6}{5}x\] Multiply both sides by 5/6: \[\frac{5}{6} \dot\ \frac{4}{3} = x\] Simplify \[\frac{5}{3} \dot\ \frac{2}{3} = x\] \[\frac{10}{9} = x\] \[1 \frac{1}{9} = x\]